ah42632:

What are the values of x and y? A. x equals 9 over 4 and y equals 3 over 4 B. x equals 9 over 4 and y equals 15 over 4 C. x equals 15 over 4 and y equals 5 over 4 D. x equals 3 over 4 and y equals 15 over 4

ah42632:

Here is the ss of the problem

KynxKatz:

Lemme see

KynxKatz:

@AZ

ah42632:

question is too hard for my brain to comprehend XD

KynxKatz:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @ah42632 question is too hard for my brain to comprehend XD \(\color{#0cbb34}{\text{End of Quote}}\) That’s y i called @AZ

AZ:

So when we have two right triangles like this there's this formula so for the triangle I'm attaching \(\sf \dfrac{BC}{AC} = \dfrac{AC}{DC}\) and \(\sf \dfrac{BC}{AB} = \dfrac{AB}{DB}\) and \(\sf \dfrac{DC}{AD} = \dfrac{AD}{DB}\)

1 attachment
AZ:

so for your triangle, let's first find 'x' and then we can just use the Pythagorean theorem to find y

AZ:

so basically if we set it up, we get \(\dfrac{4}{3} = \dfrac{3}{x}\) can you cross multiply and solve for x?

ah42632:

4x and 9?

ah42632:

or 12 and 3x?

AZ:

you got it \(\dfrac{a}{b} = \dfrac{c}{d}\) would be ad = bc so 4x = 9 what is x = ??

AZ:

just divide 4 on both sides

ah42632:

2.25

AZ:

yup!

AZ:

now the entire length of AC is 4 + 2.25 which equals what?

ah42632:

6.25

AZ:

and then since it's a big right triangle (AB)^2 + y^2 = AC^2 and AB is 5 and AC is 4 + 2.25 which is 6.25 like you said and so plug it in and solve for y

AZ:

5^2 + y^2 = 6.25^2 y = ?

ah42632:

those numbers are to the second power right?

AZ:

yes and I really have to run now so @snowflake0531 will take over :)

ah42632:

ok

snowflake0531:

Mhm, I'm going to take this a different way lol Proportions \[y/3 = 5/4\] Cross multiply

ah42632:

4y and 15

snowflake0531:

Yep~ 4y=15 Solve for y y=....

snowflake0531:

*keep it as a fraction

snowflake0531:

Divide both sides by 4

ah42632:

3 over 80?

snowflake0531:

4y=15 Divide each side by 4

ah42632:

3.75 and then as fraction I got 3/80

snowflake0531:

Division=fraction 15 divided by 4 = 15/4 So, y is 15/4 And we already know that x is 9/4 So, which one is the correct answer

ah42632:

they both are arent they? theyre both for different variables

snowflake0531:

`So, y is 15/4 And we already know that x is 9/4`

snowflake0531:

x equals 9 over 4... y equals 15 over 4

ah42632:

ohhhhhhh... I didnt read it correctly. My bad

snowflake0531:

so you got the answer?

ah42632:

yep

snowflake0531:

kay

ah42632:

thanks

snowflake0531:

yw~ altho also thank AZ xd

ah42632:

yessir @AZ thanks